# Maximum-likelihood estimation of the Matérn covariance structure of isotropic spatial random fields on finite, sampled grids

### v1

@@ 8 A C K N O W L E D G E M E N T S
This work was sponsored by the National Aeronautics and Space Administration under grant NNX11AQ45G. OLW gratefully acknowledges financial support from the Schmidt DataX Fund, grant number 22-008, at Princeton University, made possible through a major gift from the Schmidt Futures Foundation. FJS thanks the Institute for Advanced Study for a stimulating research environment during 2023–2024, and KU Leuven and the Groot Begijnhof for enabling productive Summer months. We thank Erin O’Neill for helpful comments on the final draft. We are grateful to the Associate Editor, and the anonymous reviewers, for constructive reviews that helped improve the manuscript.

### v2

@@ 8 A C K N O W L E D G E M E N T S
This work was sponsored by the National Aeronautics and Space Administration under grant NNX11AQ45G. OLW gratefully acknowledges financial support from the Schmidt DataX Fund, grant number 22-008, at Princeton University, made possible through a major gift from the Schmidt Futures Foundation. FJS thanks the Institute for Advanced Study for a stimulating research environment during 2023–2024, and KU Leuven and the Groot Begijnhof for enabling productive Summer months. We thank Erin O’Neil for helpful comments on the final draft. We are grateful to the Associate Editor Carl Tape, and to John Goff and Michael Stein for constructive reviews that improved the manuscript.